likeabossssssss:

n a grocery store shelf, there are 16 loaves of white bread and 15 loaves of wheat bread. A shopper randomly selects a loaf of bread. What is the probability it is wheat? Enter your answer to the nearest percent in the box.

likeabossssssss:

@KjSaif

Nnesha:

( number of wheat bread)/(total number of bread ) \[\huge\rm \frac{ number ~of~white~bread }{ Total ~number~of~bread }\]

Nnesha:

sorry wheat bread not white **

SkyVoltage43:

15/31

Nnesha:

\[\rm \frac{ number~of~wheat~bread }{ total~number~of~bread }=\frac{number ~of~wheat~bread}{number ~of~white~bread+number~of~wheat~bread}\]

likeabossssssss:

thx